# Towards Stable and Faithful Inpainting

### 08 Dec 2023 v1

@@ 3.2 Align MAE Prior with Stable Diffusion
Replace text-condition with MAE prior. SD is not trained on MAE priors. Fortunately, the text-condition part of SD is not utilized in inpainting, as we do not assume a text condition for inpainting task. Hence, we propose to replace the text-condition part of SD with our proposed MAE prior condition. However, as we do not fine-tune the SD, it cannot aligns well with the MAE prior. To achieve a proper guidance for SD, we introduces the alignment module to align both dimension and distribution.
@@
Dimension alignment. Particularly, we adopt the MAE prior $F_{\mathrm{MAE}}$ of size $N_{m}\times M_{m}$, where $N_{m}$ is the sequence length and $M_{m}$ is the MAE feature dimension. To align it with the SD condition of size $N_{s}\times M_{s}$, we adopt a linear layer to map the feature dimension from $M_{m}$ to $M_{s}$ and set $N_{s}=N_{m}$ to preserve the local MAE prior.
@@
Distribution alignment. After aligning the dimension, we use self-attention blocks to learn the distribution alignment for better guiding SD, leading to the condition $C_{\mathrm{MAE}}$ for SD. We train our alignment module with the same masking strategy used to train the MAE, while keeping the continuous-mask fine-tuned MAE and SD frozen. The training objective is the diffusion objective. To encourage the alignment module extract features from MAE prior, in the training stage, we use the full image to generate MAE prior instead of masked image with a chance of $p$.
@@
3.3 Align Mask-Unmask Regions during Decoding
Information loss of KL-VAE. SD performs all its generative processes in the latent space and subsequently decodes these latent codes back to image space using KL-VAE. Despite the decoder being trained to reconstruct the image, it encounters challenges associated with information loss. Particularly in tasks like inpainting, the unmasked region necessitates replacement with ground-truth values. Though Rombach et al. [39] claimed that the diffusion model should prioritize the informative semantic compression, while the VAE and GAN are used to tackle perceptual compression with high-frequency details, we argue that low-frequency semantic loss in KL-VAE could not be neglected, as verified in Figure 4 (b). Moreover, as shown in Figure 4 (a), repeated decoding with KL-VAE not only noticeably degrades high-frequency details but also shifts in semantic hue. As human is sensitive to low-frequency information changes in the image, even subtle color shifts can induce significant perceptual inconsistencies. This results in low-fidelity inpainting results, especially under the irregular mask structures.
Figure 5: Training pipeline of the decoder, enhancing mask-unmask consistency via augmentations in color and latent spaces.
@@ 3.3 Align Mask-Unmask Regions during Decoding
Figure 6: Qualitative SD1.5 inpainting results decoded by (b) vanilla KL-VAE decoder provided from [39], (c) larger decoder from [64], and (d) our augmented decoder. Our decoder can largely alleviate the mask-unmask color inconsistency.
@@
Unmask-region conditioned decoder. The aforementioned information loss is prohibitive to be addressed by the SD U-Net alone, especially given the already degraded input latent codes. Inspired by Zhu et al. [64], we design a larger decoder to preserve the global consistency in low-frequency fields. This decoder, as shown in Figure 2, involves incorporating masked images in the pixel-wise color space as additional inputs. Unlike [64], we concatenate the 0-1 mask as the extra input to more precisely locate masked regions. However, the larger decoder with even pixel-wise image inputs still fails to handle the incompatible color and texture between the original images and compressed ones in challenging scenes as verified in Figure 6 (c). The low-frequency gap between degraded and original images makes it challenging to explicitly address this issue.
Mask-unmask consistent decoder. To train the decoder to ensure the consistency between generated latent and unmasked pixels, we estimate and enlarge the inconsistency during the training of the decoder. Specifically, we propose the color and latent augmentation as shown in Figure 5 to enhance the decoder to explicitly harmonize masked regions based on unmasked pixels.
i) Color augmentation: all training images would be augmented in brightness, contrast, saturation, and hue. This color augmentation is to avoid the decoder simply learns to lighter the image, as most VAE degradation results in a darker hue and oversaturation.
ii) Latent augmentation: To simulate the real VAE degradation, we propose leveraging the artifacts generated from the diffusion model. However, denoising to real images iteratively is notably time-consuming, even with DDIM [42]. Instead, we propose the one-step estimation of degraded latent $\tilde{\bm{z}}_{0}$ via the encoded clean latent $\bm{z}_{0}$ and noised latent $\bm{z}_{t}$:
@@
$\tilde{\bm{z}}_{0}=\frac{1}{\sqrt{\overline{\alpha}_{t}}}(\hat{\bm{z}}_{t}-\sqrt{1-\overline{\alpha}_{t}}\varepsilon_{\theta}\left([\bm{z}_{t};\bm{z}_{0};\mathbf{M}],t\right)),$ (1)
@@
where the timestep $t$ is randomly sampled from $[500,1000)$; $\overline{\alpha}_{t}$ indicates the cumulative production of prescribed variance schedule in diffusion; $\varepsilon_{\theta}(\cdot)$ is the noise estimation from frozen SD based on noised $\bm{z}_{t}$, unmasked $\bm{z}_{0}$, and all-zero masking $\mathbf{M}$. Despite $\tilde{\bm{z}}_{0}$ suffering from an extremely large step denoising, we empirically find that the inpainting fine-tuned SD could produce reliable results with little artifacts enhanced by the concatenation of unmasked latent $\bm{z}_{0}$. Moreover, the latent augmentation makes the image undergo two-round encoding and decoding from KL-VAE. This not only facilitates the incorporation of natural VAE degradation into the decoder training but also enhances the overall robustness of the model. We empirically apply 50% to operate the latent augmentation to training images. We fine-tune the decoder based on [64], showcasing superior consistency as compared in Figure 6.
@@ 4 Evaluation Dataset
Figure 7: Examples of masked images constructed in MISATO. The first three cases focus on different image domains, while in the last case we focus on the inpainting of background.
To validate the inpainting performance across different domains and mask styles, we construct a evaluation dataset, dubbed as MISATO, from Matterport3D [10], Flickr-Landscape [29], MegaDept[27], and COCO 2014 [30]. Matterport3D contains a indoor image dataset, Flickr-Landscape is consisted of outdoor landscape images, MegaDepth involves a dataset of buildings, and COCO has precise segmentation of foreground subjects. For each dataset, we select 500 representative examples, yielding 2000 testing examples of size $512^{2}$.
The principle of constructing MISATO is to select the most representative and diverse examples. To this end, for Matterport3D, Flickr-Landscape and MegaDepth, we use CLIP visual model [37] to extract semantic visual features. Then we use BisectingKMeans [44] to cluster each dataset into 500 clusters, and select the cluster centers as the evaluation data. The selected data are center cropped and then resized to $512^{2}$. For COCO, we focus on the background inpainting. To this end, for each data we identify the foreground with provided segmentation and remove it from the generated masks, yielding a dataset specified for purely background inpainting.
Combined together, MISATO contains four inpainting domains, indoor, outdoor landscape, building, and background. As shown in Figure 7, for each evaluation image, we adopt the similar masking strategy as in Section 3.1 while not using the rectangle and complement rectangle masks. The masking ratio is set as $[0.2,0.8]$.
@@ 5 Experiments
Implementation details. ASUKA utilizes the stable diffusion text-guided inpainting model v1.5 as the generative backbone. We train ASUKA on Places2 [63]. For the MAE [20] used in ASUKA, we fine-tune from the MAE used in Cao et al. [8] with a batch size of 1024. We use four self-attention blocks for the alignment module, training from scratch with AdamW [32] optimizer of learning rate 5e-2 for 23K steps of batch size 64 with the standard diffusion objective. We use warm-up strategy to set initial $p$ as $100\%$ and linearly decay it to $10\%$ in the first 2k training steps and then freeze $p=10\%$. We fine-tune the decoder based on [64] for 50K steps with a batch size of 40 and learning rate of 8e-5 with cosine decay. We use ColorJitter to implement color augmentation, with the setting of brightness 0.15, contrast 0.2, saturation 0.1, and hue 0.03.
Evaluation metrics. We compare inpainting algorithms using the following metrics: Peak Signal-to-Noise Ratio (PSNR) to compare the ratio of real data to introduced noise, Structural Similarity Index Measure (SSIM) [47] to identify the perceived change in structural information, Learned Perceptual Image Patch Similarity (LPIPS) [60] to calculate the patch-level image distances, Fréchet Inception Distance (FID) [21] to compare the distribution distance between generated images and real images, and Paired/Unpaired Inception Discriminative Score (P-IDS/U-IDS) [61] to measure the human-inspired linear separability.
5.1 Comparison on MISATO
Figure 8: Qualitative comparison on MISATO. We compare ASUKA with Co-Mod [61], MAT [26], LaMa [45], MAE-FAR [8], and SD [39] in landscape (first two rows), indoor (3rd and 4th rows), building (5th and 6th rows) and background inpaintings (last two rows). SD performs unconditional generation. SD-text utilizes text prompt of “background”. SD-prompt utilizes trained prompt for inpainting task using the same training setting of ASUKA. Inpainting algorithms usually suffer from fidelity issue, while SD-based algorithms suffer from stability issue and generate random elements. In contrast, ASUKA achieves stable and faithful inpainting.
@@ 5.1 Comparison on MISATO
Co-Mod [61] 21.805 0.787 0.179 17.421 0.243 0.109
MAT [26] 21.553 0.787 0.176 17.261 0.255 0.122
LaMa [45] 23.677 0.820 0.155 15.436 0.260 0.135
MAE-FAR [8] 24.150 0.821 0.142 13.283 0.282 0.153
SD [39] 22.454 0.796 0.168 12.812 0.345 0.211
SD-text 22.394 0.793 0.164 12.603 0.337 0.207
SD-prompt 22.381 0.798 0.160 12.517 0.331 0.204
ASUKA 22.495 0.799 0.150 11.460 0.368 0.256
Table 1: Quantitative comparison on MISATO. Co-Mod [61], MAT [26], LaMa [45], and MAE-FAR [8] are state-of-the-art inpainting algorithms. SD [39] performs unconditional generation. SD-text utilizes text prompts of “background” to perform text-guided generation. SD-prompt utilizes learnable prompts trained specifically for inpainting, using the same training setting as ASUKA, performing prompt-guided generation. ASUKA and SD use the stable diffusion text-guided inpainting model v1.5.
Competitors. We first compare ASUKA with SD [39]. Specifically, we consider three variants of SD: i) SD, which generates inpainting results with no guidance, and performs unconditional generation; ii) SD-text, which generates inpainting results guided by the text prompt of “background”; iii) SD-prompt, where we replace the text prompt with learnable tokens and train the tokens using the same training pipeline of ASUKA. Beyond SD, we also compare ASUKA with state-of-the-art inpainting algorithms, including Co-Mod [61], MAT [26], LaMa [45], and MAE-FAR [8].
Quantitative comparison results are reported in Table 1. ASUKA, although based on frozen SD, achieves significant and consistent improvements over SD across all the evaluation metrics, and enjoys the state-of-the-art results on FID, U-IDS and P-IDS. Particularly, the U-IDS and P-IDS coincide with human preference [61], and has a potential upper bound of $0.5$, indicating the superiority of ASUKA by a large margin. The inconsistency results between PSNR, SSIM, LPIPS and FID, U-IDS, P-IDS, partially comes from the frozen SD, where ASUKA achieves consistent improvements but is bounded to the frozen U-Net. On the other hand, the PSNR and SSIM is known to have bad correlation with human preferences, as several illustrative examples shown in the appendix. In summary, the results validate the contribution of ASUKA that yields stable and faithful inpainting through frozen SD.
Qualitative comparison examples are shown in Figure 8. i) The state-of-the-art inpainting algorithms usually suffer from fidelity issues, for example the unnatural boundaries in the third and fourth rows, and failed inpainting of tower in the third-to-last row. LaMa and MAE-FAR sometimes lead to blurred inpainting results, especially in the scenario of large continuous masks. ii) The three variants of SD usually suffer from the stability issue and generate random new elements in the masked region, in almost all the illustrated images. iii) Particularly, for the background mask shown in the second-to-last row, state-of-the-art inpainting algorithms cannot identify the object precisely, and extend the human in hairs and elbows. On the other hand, SD generates random elements in the masked region. iv) In contrast, ASUKA enjoys superior stable and faithful inpainting.

### 17 Mar 2024 v2

@@ 3.2 Align MAE Prior with SD for Context-Stability
Replace text-condition with MAE prior. SD is not trained on MAE priors. Fortunately, the text-condition part of SD is not utilized in inpainting, as we do not assume a text condition for inpainting task. Hence, we propose to replace the text-condition part of SD with our proposed MAE prior condition. However, as we do not fine-tune the SD, it cannot aligns well with the MAE prior. To achieve a proper guidance for SD, we introduce the alignment module to align MAE with SD in both dimension and distribution perspective, as shown in 3(b).
@@
Dimension alignment. Particularly, the MAE prior $F_{\mathrm{MAE}}$ is of size $N_{m}\times M_{m}$, where $N_{m}$ is the sequence length and $M_{m}$ is the feature dimension. To align it with the SD condition of size $N_{s}\times M_{s}$, we adopt a linear layer to map the feature dimension from $M_{m}$ to $M_{s}$ and set $N_{s}=N_{m}$ to preserve the local MAE prior.
@@
Distribution alignment. After aligning the dimension, we use self-attention blocks to learn to better guiding SD, leading to the condition $C_{\mathrm{MAE}}$ for SD. We train our alignment module using the standard diffusion objective with the same masking strategy used to train the MAE, keeping other modules frozen.
Misalignment. When training the alignment module with the set (input image, MAE prior, SD result), misalignment issues may arise. For example, if an object is completely masked, the MAE will predict the masked area with background, whereas the SD is trained to recreate the object. This difference can lead the alignment module to mistakenly disregard the MAE prior. To address this, we improve the SD’s adherence to the MAE prior by substituting the MAE predicted prior with the MAE reconstructed prior at a probability of $p$. The MAE reconstructed prior involves using MAE to recreate the image without masking any area, ensuring MAE has access to all information needed for reconstruction. This approach helps train the alignment module to better guide the SD.
@@
(a) Qualitative SD1.5 inpainting results decoded by (b) vanilla decoder of SD [39], (c) larger decoder [62], (d) our decoder. Our decoder largely alleviate the mask-unmask color inconsistency.
(b) Decoder trained by local harmonization task, enhancing mask-unmask consistency by reconstructing original image guided by the unmasked region from augments in color and latent spaces.
Figure 6: ASUKA trains a specialized-decoder to solve the color-inconsistency issue from the information loss of KL-VAE and domain gap between generated and real latent.
3.3 Mask-Unmask Align during Decoding for Visual-Consistency
Information loss of KL-VAE. SD performs all its generative processes in the latent space and subsequently decodes these latent codes back to image space using KL-VAE. Despite the decoder being trained to reconstruct the image, it encounters challenges associated with information loss. Particularly in tasks like inpainting, we have ground-truth values for the unmasked region. Though Rombach et al. [39] claimed that the diffusion model should prioritize the informative semantic compression, while the VAE and GAN are used to tackle perceptual compression with high-frequency details, we argue that low-frequency semantic loss in KL-VAE could not be neglected, as verified in Fig. 5 (b). The KL-VAE will not only noticeably degrades high-frequency details but also shifts in semantic hue. This shift can be verified by repeated reconstruction with KL-VAE, as shown in Fig. 5 (a) where larger shift is observed during repeated reconstruction. As human is sensitive to low-frequency information changes in the image, even subtle color shifts can induce significant visual-inconsistencies. This visual-inconsistency issue is more severe under the irregular or large mask structures.
@@ 3.3 Mask-Unmask Align during Decoding for Visual-Consistency
Unmask-region conditioned decoder. The aforementioned information loss is prohibitive to be addressed by the SD U-Net alone, especially given the already degraded input latent codes. Zhu et al. [62] adopts a larger decoder, with additional inputs of masked images. However, it still fails to handle the incompatible color and texture between the original images and compressed ones in challenging scenes as verified in Fig. 5(a) (c). The low-frequency gap between degraded and original images makes it challenging to explicitly address this issue. To solve the mask-unmask visual-inconsistency issue, we reformulate the decoding as a local harmonization task to preserve the visual consistency. As shown in Fig. 3, compared with SD decoder, our decoder involves additional inputs of masked images in the pixel-wise color space and the 0-1 mask.
Mask-unmask visual-consistent decoder. To train the decoder to ensure visual-consistency between generated latent and unmasked pixels, we re-formulate the decoding as a local harmonization task. Specifically, we propose the color and latent augmentation as shown in Fig. 5(b) to estimate and enlarge the visual-inconsistency during decoding. We follow the standard KL-VAE training pipeline of SD, but replacing the inputs with augmented ones. Particularly, we use the original image as the reconstruction target and use color and latent augmentation to corrupt input image, simulating the information loss of KL-VAE and domain gap between generated and real latent, respectively. This forces the decoder to reconstruct the un-corrupted image based on the ground-truth unmasked region.
i) Color augmentation: The first information loss comes from the imperfect VAE reconstruction. VAE is trained to encode and decode the input image for reconstruction, but suffers from color shift as shown in Fig. 5(a) (b). Empirically, further conditioned on unmasked image alleviate but not solve the color inconsistency issue, as shown in Fig. 5(a) (c). Hence, we need to explicitly train the decoder to ensure color consistency. To this end, we augment all training images in brightness, contrast, saturation, and hue, and requires the decoder to reconstruct original image conditioned on the unaugmented unmasked image. This encourages the decoder to fully utilize the information from unmasked regions.
ii) Latent augmentation: The other information loss comes from the distribution gap between SD generated latent and real latent. This gap also causes color inconsistency even if we alleviate the VAE reconstruction loss, as shown in Fig. 6(a). To simulate this loss, we incorporate the artifacts generated from the SD to train the decoder. However, denoising to real images iteratively is notably time-consuming, even with DDIM [42]. To balance the efficiency and efficacy, we design a one-step estimation of SD. As our target is to capture the generation gap, we use the clean latent $\bm{z}_{0}$ and all-zero mask $\mathbf{M}$ as conditions. This tells the SD all the needed information to generate the clean latent, ensuring the generated latent preserves content and only shift from the distribution gap. We follow the standard pipeline to estimate $\bm{z}_{0}$ with modified conditions as:
@@
$\hat{\bm{z}}_{0}=\frac{1}{\sqrt{\overline{\alpha}_{t}}}(\bm{z}_{t}-\sqrt{1-\overline{\alpha}_{t}}\varepsilon_{\theta}\left([\bm{z}_{t};\bm{z}_{0};\mathbf{M}],t\right)),$ (1)
@@
where the timestep $t$ is randomly sampled from $[500,1000)$; $\overline{\alpha}_{t}$ indicates the prescribed variance schedule in SD; $\varepsilon_{\theta}(\cdot)$ is the frozen SD take as inputs noised $\bm{z}_{t}$, unmasked $\bm{z}_{0}$, and all-zero masking $\mathbf{M}$. The large step denoising is chosen to increase the distribution gap, as empirically SD could produce reliable results in small $t$ given the unmasked latent condition $\bm{z}_{0}$. Then we decode $\hat{\bm{z}}_{0}$ to image as the latent augmented inputs. This makes latent augmentation an off-line strategy. We apply latent augmentation to 50% training images. We fine-tune the decoder based on [62], showcasing superior consistency as compared in Fig. 5(a).
(a) w/ v.s. w/o latent aug.
(b) Different image domains in MISATO.
Figure 7: (a) Ablation of latent augmentation; (b) MISATO evaluation examples.
@@ 4 Evaluation Dataset
[…]
In addition to the standard benchmark Places 2 [61], to validate the inpainting performance across different domains and mask styles, we construct a evaluation dataset, dubbed as MISATO, from Matterport3D [10], Flickr-Landscape [28], MegaDept[26], and COCO 2014 [29]. The first three datasets contains images from indoor, outdoor landscape and building, respectively. COCO has object segmentation. We select 500 representative examples of size $512^{2}$ from each dataset. The MISATO dataset is available at https://github.com/Yikai-Wang/asuka-misato.
The principle of constructing MISATO is to select the most representative and diverse examples. To this end, for first three datasets, we use CLIP visual model [37] to extract semantic visual features. Then we use BisectingKMeans [44] to cluster each dataset into 500 clusters, and select the cluster centers as the evaluation data. The selected data are center cropped and then resized to $512^{2}$. For COCO, we focus on the background inpainting. To this end, for each data we identify the foreground with provided segmentation and remove it from the generated masks, yielding a dataset specified for purely background inpainting.
Combined together, MISATO contains 2000 examples from four inpainting domains, indoor, outdoor landscape, building, and background. As shown in Fig. 6(b), we adopt the masking strategy as in Sec. 3.1 but excluding the rectangle and complement rectangle masks. The masking ratio is set as $[0.2,0.8]$.
@@ 5 Experiments
Implementation details. ASUKA utilizes the stable diffusion text-guided inpainting model v1.5 as the generative backbone. We train ASUKA on Places2 [61]. For the MAE [19] used in ASUKA, we train on images of size $256^{2}$, which is efficient and produce context-stable guidance for SD to generate high-resolution images. We fine-tune from the pre-trained MAE with a batch size of 1024. We use four self-attention blocks for the alignment module, training from scratch with AdamW [31] of learning rate 5e-2 with the standard diffusion objective. We set $p$ as $100\%$ and linearly decay it to $10\%$ in the first 2K training steps and then freeze. We fine-tune the decoder based on [62] for 50K steps with a batch size of 40 and learning rate of 8e-5 with cosine decay. We use ColorJitter for color augmentation, with brightness 0.15, contrast 0.2, saturation 0.1, and hue 0.03.
Evaluation metrics. We use the Learned Perceptual Image Patch Similarity (LPIPS) [58] to calculate the patch-level image distances, Fréchet Inception Distance (FID) [20] to compare the distribution distance between generated images and real images, and Paired/Unpaired Inception Discriminative Score (P-IDS/U-IDS) [59] to measure the human-inspired linear separability.
Competitors. We consider three variants of SD [39]: i) SD, prompted by "", performing unconditional generation; ii) SD-text, prompted by "background" as we have no captions in inpainting; iii) SD-prompt, prompted by learnable tokens which are trained using the same training pipeline of ASUKA. We also compare with state-of-the-art inpainting algorithms Co-Mod [59], MAT [25], LaMa [45], MAE-FAR [7], and SD-Repaint [32].
Table 1: Quantitative comparison on MISATO and Places 2. Co-Mod [59], MAT [25], LaMa [45], MAE-FAR [7] and SD-Repaint [32] are state-of-the-art inpainting methods. SD [39] performs unconditional generation. SD-text uses “background” text prompt to guide generation. SD-prompt uses learnable prompts trained specifically for inpainting, using the same training setting as ASUKA, performing prompt-guided generation. ASUKA and SD variants use the stable diffusion text-guided inpainting model v1.5.
Dataset MISATO (2k images) Places 2 (36.5k images)
@@
Co-Mod [59] 0.179 17.421 0.243 0.109 0.267 5.794 0.274 0.096
MAT [25] 0.176 17.261 0.255 0.122 0.202 3.765 0.348 0.195
LaMa [45] 0.155 15.436 0.260 0.135 0.202 6.693 0.247 0.050
MAE-FAR [7] 0.142 13.283 0.282 0.153 0.174 3.559 0.307 0.105
SD [39] 0.168 12.812 0.345 0.211 0.193 1.514 0.375 0.207
SD-text 0.164 12.603 0.337 0.207 0.191 1.506 0.373 0.202
SD-prompt 0.160 12.517 0.331 0.204 0.189 1.477 0.390 0.234
SD-Repaint [32] 0.227 27.861 0.016 0.007 0.251 12.466 0.217 0.045
ASUKA 0.150 11.460 0.368 0.256 0.183 1.230 0.413 0.287
5.1 Comparison on Benchmarks
Quantitative comparison results are reported in Tab. 1. ASUKA, although based on frozen SD, achieves significant and consistent improvements over SD across all the evaluation metrics, and enjoys the state-of-the-art results on FID, U-IDS and P-IDS. Particularly, the U-IDS and P-IDS coincide with human preference [59], and has a potential upper bound of $0.5$, indicating the superiority of ASUKA by a large margin. The inconsistency results between LPIPS and FID, U-IDS, P-IDS, partially comes from the frozen SD, where ASUKA achieves consistent improvements but is bounded to the frozen U-Net. In summary, the results validate the contribution of ASUKA that yields context-stable and visual-consistent inpainting through frozen SD. This improvement is both validated in in-distribution dataset Places 2 and out-of-distribution dataset MISATO.
Qualitative comparison examples are shown in Fig. 8. i) The state-of-the-art inpainting algorithms usually suffer from unnatural generation, for example the unnatural boundaries in the third and fourth rows, and failed inpainting of tower in the third-to-last row. LaMa and MAE-FAR sometimes lead to blurred inpainting results, especially in the scenario of large continuous masks. ii) The SD variants usually suffer from the context-stability issue and hallucinate unreasonable objects, in almost all the illustrated images. iii) Particularly, for the background mask shown in the second-to-last row, state-of-the-art inpainting algorithms cannot identify the object precisely, and extend the human in hairs and elbows. SD hallucinates random elements in the masked region. iv) In contrast, ASUKA enjoys context-stable and visual-consistent inpainting.